OpenStudy (anonymous):

Sandy owns a restaurant. She charges $5.50 for 3 eggs and 2 pieces of toast, and $3.50 for 2 eggs and 1 piece of toast. a. How much does Sandy charge per egg? b. How much does Sandy charge per piece of toast?

OpenStudy (misty1212):

\[3\times E+2\times T=5.50\] is one equation

OpenStudy (anonymous):

I'm not really sure what you mean, can you explain?

OpenStudy (misty1212):

that one says 3 eggs and two pieces of toast cost $5.50

OpenStudy (anonymous):

Yes it does

OpenStudy (misty1212):

which one says 2 eggs and one piece of toast cost $3.50?